Sol 8 Summary – April 30th

Crew 193 Sol Summary Report – 30 APR 2018

Sol 8

Summary Title: “It was a good Sol” – Ice Cube

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Nominal

Sol Activity Summary: The day started off later than normal as the first EVA was not scheduled to leave until 0800. EVA 14 was able to explore the White Moon area on Mars! While EVA 14 was occurring, the two crew members in the hab took care of the plants in the green Hab and cleaned the shared living area in the main Hab. One crew member credits the growth of the plants to the time she spends talking to them. Maybe Mark Watney should have tried that!

The crew lunched together (thanks again to our commander for cooking), had a short meeting to continue team development and then it was time to get ready for EVA 15. EVA 15 was cut short at 1800 due to approaching rain clouds. The EVA was in the Candor Chasma area and it was decided that if it rained it would be a very difficult area to exit and return to the Hab.

Look Ahead Plan: EVA 16 will be exploring Lith Canyon and the Dinosaur Quarry. EVA 17 will explore the area around Zubrin’s Head and Southwest. We will continue to work on learning ASL for 5 minutes each day as well as sharing stories about our lives so that the crew gets to know each other better. Our daily team meetings have been working well and will be continued through the remainder of the mission.

Anomalies in work:

1. None to report today

Weather: Partly cloudy in AM, cloudy and gusty in PM with rain in the distance.

Crew Physical Status: Minor physical ailments (blister, etc.) but otherwise nominal!


EVA 14 – Explored the moon area. Continued the GPS study and improved on their ATV skills by taking their most challenging road off of Cowdung Road yet.

EVA 15 – Successfully explored the Candor Chasma area and obtained some spectacular photographs. EVA was cut short early due to the possibility of incoming rain making it difficult to exit the Candor Chasma area.

**Please note that the EVA numbers submitted in the EVA report are not correct. It should have been EVA 14 and 15, not 15 and 16**

Reports filed:


Science Report

Astronomy Report

Operations Report

EVA Report

Reports to be filed:

Commander’s Report

Photo & Photo of the Day Report

Next Day EVA Request

Support Requested:

1. Hot Tub? (No requests at this time)

Copyright © The Mars Society. All rights reserved. | Main Site